OpenStudy (anonymous):

The surface area of a cylinder with height, h, and base radius, r, is given by A = 2 rh + 2 r2. Using 3.14 for , find the surface area of a cylinder with radius of 4 cm and a height of 20 cm.

OpenStudy (radar):

Wheres the pie?

OpenStudy (radar):

\[A =2\pi rh +2\pi r ^{2}\] Isn't that the formula?

OpenStudy (radar):

A = (2)(3.14)(4)(20) + (2)(3.14)(4)(4)

OpenStudy (radar):

You can take it from there.

OpenStudy (anonymous):

Thank you for your help! I was a little lost with this problem.

OpenStudy (radar):

No problem and good luck with your studies.